# L4 collaboration, Leipzig-Lorraine-Lviv-Coventry, Europe

### v1

@@ I Introduction
Leading singularities of thermodynamic quantities exhibiting critical behavior at a second-order phase transition are characterized by power laws in the reduced temperature $t$ and external field $h$. In the thermodynamic limit, at $h=0$, the correlation length behaves as $\xi_{\infty}(t)\sim|t|^{-\nu}$, and the specific heat, susceptibility, and low-temperature magnetization respectively as $C_{\infty}(t)\sim|t|^{-\alpha}$, $\chi_{\infty}(t)\sim|t|^{-\gamma}$, and $m_{\infty}(t)\sim|t|^{\beta}$. The exponents that appear in these power laws define a universality class and identical behaviors are expected for different systems in the same universality class, usually determined by general properties like ground-state symmetries, space dimensionality, and interaction range.
@@
$\chi_{\infty}(t)\sim|t|^{-\gamma}|\ln{|t||^{\hat{\gamma}}}\;\;;\;\;\;\;m_{\infty}(t)\sim|t|^{\beta}|\ln{|t||^{\hat{\beta}}},$
@@
The presence of logarithmic corrections has already been studied numerically through Lee-Yang zeros, first by Kenna and Lang for the four-dimensional (4D) Ising model [3] and from thermodynamic quantities by Luijten [4], while then revisited by Lorenzo using numerical simulations and very large system sizes [5]. Below the upper critical dimension logarithmic corrections appear in certain models as well. For example the Lee-Yang zeros of the 2D XY model have been studied by Kenna and Irving [6, 7], as well as from Hong [8] employing tensor network calculations, but also from standard thermodynamic observables, such as the magnetic susceptibility [9, 10].
@@
Building on the preliminary results presented in Ref. [11], we propose a method to investigate the phase diagram of the 3D Blume-Capel model. Our approach aims to refine the transition line by analyzing the finite-size scaling behavior of Lee-Yang zeros and high-order cumulants of the magnetization, while taking into account corrections to scaling. To the best of our knowledge, there has been no dedicated numerical analysis of the tricritical exponents in the Blume-Capel model that incorporates logarithmic corrections. The availability of both analytical and highly accurate numerical values for the leading exponents, as well as for those related to the corrections-to-scaling at the tricritical point and along the second-order transition line (Ising universality), enables us to impose stringent constraints on the phase diagram. In particular we study the model at (i) its zero crystal-field critical point and (ii) exactly at the tricritical point through hybrid Monte-Carlo simulations that combine in an optimized manner the Wolff cluster update with the Metropolis single spin-flip dynamics. For the determination of the first zero we use the cumulant method, originally developed by Deger et al., and successfully tested in various models [16, 17, 18, 19, 20].
@@ II.2 Numerical details
We employ a hybrid numerical scheme that combines effectively a Wolff single-cluster update [34] for the ±$\pm 1$ spins and a single-spin-flip Metropolis update [35, 36, 25] to account for the vacancies. Our simulations are enhanced by a histogram reweighting method, which allows us to extrapolate data obtained from simulations at fixed values of the crystal field (temperature) to nearby temperature (crystal-field) ranges. Using this hybrid scheme we simulate, under periodic boundary conditions, systems with linear sizes in the range $L=12-22$ (hereafter, $N=L\times L\times L$ defines the total number of spins on the lattice). Following the previously successfully tested prescription of Refs. [37, 23], in this hybrid approach an elementary Monte Carlo step is defined as $3N$ Metropolis steps followed by $L$ Wolff steps. We perform $900\times N$ Monte Carlo steps per spin to ensure equilibration, followed by $900\times 5\times N$ Monte Carlo steps per spin for the collection of numerical data. As already mentioned above, our simulations are carried out at the two special points along the phase boundary, the zero crystal field and the tricritical point using the literature results $(\Delta_{\rm c},T_{\rm c})=(0,3.1952)$ [26] and $(\Delta_{\rm t},T_{\rm t})=(2.8450,1.4182)$ [28], respectively. A comment for the application of finite-size scaling to the numerical data: We restrict ourselves to data with $L\geq L_{\rm min}$. As usual, to determine an acceptable $L_{\rm min}$ we make use of the standard $\chi^{2}$ test of goodness of fit [38]. Finally, errors are computed using the Jackknife method [39].
@@ II.3 Observables
To extract the critical behavior of the model we focus on the usual thermodynamic quantities, such as the magnetic susceptibility
@@
$\chi_{2}=\frac{\langle E_{\Delta}^{2}\rangle-\langle|E_{\Delta}|\rangle^{2}}{k_{\rm B}T},$ (5)
as well as $\chi_{12}$, a quantity similar to the magnetocaloric coefficient, obtained exclusively from the crystal-field contribution to the energy
@@
Figure 1: Finite-size scaling behavior of: (a) the Lee-Yang zeros at criticality from the fit $h_{0}\sim aL^{-y_{h}}$ and (b) the magnetization cumulant at criticality from $\langle\langle M^{8}\rangle\rangle\sim aL^{8y_{h}-3}$, both for a set of temperatures around $T=3.1962$. Note the value of the exponent $y_{h}$ marked on the left vertical scale, and the quality of the fit $\chi^{2}_{s}$ on the right vertical scale.
An alternative approach to studying critical phenomena involves analyzing specific cumulants of thermodynamic observables and the zeros of the partition function (since these are associated with critical singularities in the free energy). In fact as it was shown explicitly in Ref. [20] for the $d=2$ Blume-Capel model, the behavior of the magnetization cumulants is designed for an accurate determination of the underlying critical exponents even from the study of very small system sizes. The nth moment of the magnetization (here in the presence of a magnetic field) can be defined as the nth derivative of the free energy with respect to that magnetic field
@@
whose finite-size scaling follows the ansatz $\langle\langle M^{n}(0)\rangle\rangle\sim L^{-d+ny_{h}}$. Returning now to the case of Lee-Yang zeros, which are the remaining quantities of interest, these can be also defined from the magnetization cumulants via [16, 17, 18, 19]
@@ III.1 Critical Ising scaling behavior
Figure 2: Finite-size scaling analysis at criticality of the first Lee-Yang zero from fits of the form $h_{0}L^{1.5874}\sim a(1+bL^{-0.832}+cL^{-1.168})$ [panel (a)] and $h_{0}L^{1.5874}\sim a(1+bL^{-0.832})$ panel (b)]. Similar analysis for the $8^{\rm th}$-order magnetization cumulant using now fits of the form $\langle\langle M^{8}\rangle\rangle L^{-16.8552}\sim a(1+bL^{-0.832}+cL^{-1.168})$ [panel (c)] and $\langle\langle M^{8}\rangle\rangle L^{-16.8552}\sim a(1+bL^{-0.832})$ [panel (d)].
@@ III.1.2 Accurate determination of the critical temperature
Figure 3: Finite-size scaling analysis of the density of the first Lee-Yang zero close to criticality for three values of the temperature, as indicated.
@@
In Fig. 3, we present the results for three different temperatures, as indicated in the panel. The curves are nearly indistinguishable, yet the best fit quality is achieved for the data corresponding to $T_{\rm c}=3.1962$, further reinforcing the findings from the previous analysis based on the first Lee-Yang zero and the magnetization cumulant.
@@
Additionally, the measured value for the exponent, $(d/y_{h})\approx 1.2057$ is in excellent agreement with the expected value, showing a deviation of less than $1\%$.
To move forward, it is now instructive to analyze more traditional quantities, namely the standard thermodynamic quantities. This will not only help us verify whether the results provide a consistent picture but also allow us to assess the advantages of our current approach. An estimate of the critical temperature is typically found by studying the crossings of the Binder cumulant (or of some other universal ratio, like $\xi/L$, where $\xi$ the correlation length) or via the shift behavior of related pseudocritical temperatures measured for the various diverging quantities (for example the location of the peaks of the specific heat or susceptibility). We choose to examine this latter route here for a set of thermodynamic observables defined above in Sec. II.3. As one also expects corrections-to-scaling, the scaling law describing the shift behavior of these pseudocritical temperatures $T_{L}$ reads as
@@ III.2 Tricritical scaling behavior
As previous estimations in the literature place the tricritical point near $(\Delta_{\rm t},T_{\rm t})=(2.8446,1.4182)$, we shall explore several values of these parameters in the vicinity of this point. The finite-size scaling analysis at both estimates proposed in Refs. [28, 27] for the location of the tricritical point gives a positive slope for the curves of the Lee-Yang zeros – $\ln(h_{0}L^{5/2})$ versus $\ln(\ln L)$ – where a negative slope is expected due to the $a(\ln L)^{-1/6}$ prediction in Eq. (22). A similar discrepancy appears for the nth-order magnetization cumulant for which a negative slope is measured, whereas one expects a positive slope, namely $\sim(\ln L)^{n/6}$ in Eq. (23). We therefore focus on three neighbouring temperatures: $T=1.4181$, $1.4182$, and $T=1.4183$. Additionally, we examine four nearby values of the crystal field within the range $\Delta=2.8441-2.8444$. The numerical data in Fig. 6 indicate that the exponents are highly sensitive to the value of the crystal field, with variations observed even up to the fourth decimal place. Knowing that the expected exponent is $\hat{y}_{h}^{\rm(ref)}=1/6$, we find compatible results when $\Delta=2.8442$, along with a very good quality of fit. Indeed, at $T=1.4181$, the exponent value and the fit quality are slightly less accurate compared to those at $T=1.4182$ and $T=1.4183$. For these latter temperatures we obtain $\hat{y}_{h}=0.172(8)$ with $\chi^{2}_{s}=1.68$ and $\hat{y}_{h}=0.158(8)$ with $\chi^{2}_{s}=1.35$.
@@ Acknowledgements.
This work began in collaboration with our dear friend Ralph Kenna, who sadly passed away a year ago. In his memory, we would like to highlight two special issues published in tribute to him: one in the Ukrainian journal Condensed Matter Physics [51] published by the ICMP, which Ralph had strong and fruitful collaborations with, and another in the journal Entropy [52], where he served on the editorial board. L.M. would like to express her gratitude to Ralph who was her PhD co-supervisor, for equipping her with the essential tools to advance her work. L.M. also thanks Coventry University, the L4 collaboration Leipzig-Lorraine-Lviv-Coventry and the UFA (French-German University) through the Collège Doctoral franco-allemand 02-07 for their support, and Denis Gessert for enriching discussions. N.G.F. would like to dedicate this paper to the memory of Ralph Kenna, an inspiring scientist and close colleague for more than 10 years at Coventry University. The work of N.G.F. was supported by the Engineering and Physical Sciences Research Council (grant EP/X026116/1 is acknowledged).

### v2

@@ I Introduction
Leading singularities of thermodynamic quantities exhibiting critical behavior at a second-order phase transition are characterized by power laws in the reduced temperature $t$ and external field $h$. In the thermodynamic limit, at $h=0$, the correlation length behaves as $\xi_{\infty}(t)\sim|t|^{-\nu}$, and the specific heat, susceptibility, and low-temperature magnetization respectively as $C_{\infty}(t)\sim|t|^{-\alpha}$, $\chi_{\infty}(t)\sim|t|^{-\gamma}$, and $m_{\infty}(t)\sim|t|^{\beta}$ (for $t<0$). The exponents that appear in these power laws define a universality class and identical behaviors are expected for different systems in the same universality class, usually determined by general properties like ground-state symmetries, space dimensionality, and interaction range.
@@
$\chi_{\infty}(t)\sim|t|^{-\gamma}|\ln{|t||^{\hat{\gamma}}}\;\;;\;\;\;\;m_{\infty}(t)\sim|t|^{\beta}|\ln{|t||^{\hat{\beta}}}\text{ for $t<0$},$
@@
The presence of logarithmic corrections has already been studied numerically through Lee-Yang zeros, first by Kenna and Lang for the four-dimensional (4D) Ising model [3] and from thermodynamic quantities by Luijten [4], while then revisited by Ruiz-Lorenzo using numerical simulations and very large system sizes [5]. Below the upper critical dimension logarithmic corrections appear in certain models as well. For example the Lee-Yang zeros of the 2D XY model have been studied by Kenna and Irving [6, 7], as well as from Hong [8] employing tensor network calculations, but also from standard thermodynamic observables, such as the magnetic susceptibility by Janke  [9] and Hasenbusch [10].
@@
Building on the preliminary results presented in Ref. [11], we propose a method to investigate the phase diagram of the 3D Blume-Capel model. Our approach aims to refine the transition line by analyzing the finite-size scaling behavior of Lee-Yang zeros and high-order cumulants of the magnetization, while taking into account corrections to scaling. To the best of our knowledge, there has been no dedicated numerical analysis of the tricritical exponents in the Blume-Capel model that incorporates logarithmic corrections. The availability of both analytical and highly accurate numerical values for the leading exponents, as well as for those related to the corrections-to-scaling at the tricritical point and along the second-order transition line (Ising universality class), enables us to impose stringent constraints on the phase diagram. In particular we study the model at (i) its zero crystal-field critical point and (ii) exactly at the tricritical point through hybrid Monte-Carlo simulations that combine in an optimized manner the Wolff cluster update with the Metropolis single spin-flip dynamics. For the determination of the first zero we use the cumulant method, originally developed by Deger et al., and successfully tested in various models [16, 17, 18, 19, 20].
@@ II.2 Numerical details
We employ a hybrid numerical scheme that combines effectively a Wolff single-cluster update [34] for the ±$\pm 1$ spins and a single-spin-flip Metropolis update [35, 36, 25] to account for the vacancies. Our simulations are enhanced by a histogram reweighting method, which allows us to extrapolate data obtained from simulations at fixed values of the crystal field (temperature) to nearby temperature (crystal-field) ranges. Using this hybrid scheme we simulate, under periodic boundary conditions, systems with linear sizes in the range $L=12-22$ (hereafter, $N=L\times L\times L$ defines the total number of spins on the lattice). Following the previously successfully tested prescription of Refs. [37, 23], in this hybrid approach an elementary Monte Carlo step is defined as $L$ Wolff steps and $3N$ Metropolis steps: more precisely, after each Wolff step, $3L$ Metropolis spin flips are carried out. We perform $900\times N$ Monte Carlo steps per spin to ensure equilibration, followed by $900\times 5\times N$ Monte Carlo steps per spin for the collection of numerical data. As already mentioned above, our simulations are carried out at the two special points along the phase boundary, the zero crystal field and the tricritical point using the literature results $(\Delta_{\rm c},T_{\rm c})=(0,3.1952)$ [26] and $(\Delta_{\rm t},T_{\rm t})=(2.8450,1.4182)$ [28], respectively. A comment for the application of finite-size scaling to the numerical data: We restrict ourselves to data with $L\geq L_{\rm min}$. As usual, to determine an acceptable $L_{\rm min}$ we make use of the standard $\chi^{2}$ test of goodness of fit [38]. Finally, errors are computed using the Jackknife method [39].
@@ II.3 Observables
To extract the critical behavior of the model we focus on the usual thermodynamic quantities, the thermal average are denoted as $\langle O\rangle$ where $O$ is the thermodynamic quantity. The magnetic susceptibility is written
@@
$\chi_{2}=\frac{\langle E_{\Delta}^{2}\rangle-\langle E_{\Delta}\rangle^{2}}{k_{\rm B}T},$ (5)
where $E_{\Delta}=\sum_{i}\sigma_{i}^{2}$ is the crystal field energy. One can also define $\chi_{12}$, a quantity similar to the magnetocaloric coefficient, obtained exclusively from the crystal-field contribution to the energy
@@
Figure 1: Finite-size scaling behavior of: (a) the first Lee-Yang zero at criticality from the fit $h_{0}\sim aL^{-y_{h}}$ and (b) the magnetization cumulant at criticality from $\langle\langle M^{8}\rangle\rangle\sim aL^{8y_{h}-3}$, both for a set of temperatures around $T=3.1962$. Note the value of the exponent $y_{h}$ marked on the left vertical scale, and the quality of the fit $\chi^{2}_{s}$ on the right vertical scale.
An alternative approach to studying critical phenomena involves analyzing specific cumulants of thermodynamic observables and the zeros of the partition function (since these are associated with critical singularities in the free energy). In fact as it was shown explicitly in Ref. [20] for the $d=2$ Blume-Capel model, the behavior of the magnetization cumulants $\langle\langle M^{n}\rangle\rangle$ (the double brackets refer to the cumulants) is designed for an accurate determination of the underlying critical exponents even from the study of very small system sizes. The nth moment of the magnetization (here in the presence of a magnetic field) can be defined as the nth derivative of the free energy with respect to that magnetic field
@@
whose finite-size scaling follows the Ansätze $\langle\langle M^{n}(0)\rangle\rangle\sim L^{-d+ny_{h}}$ where $y_{h}$ is the renormalization-group eigenvalue associated to the magnetic field $h$, also defined as $y_{h}=\beta\delta/\nu$. Returning now to the case of Lee-Yang zeros, which are the remaining quantities of interest, the first zero can be defined from the magnetization cumulants for large $n$ via [16, 17, 18, 19]
@@ III.1 Critical Ising scaling behavior
Figure 2: Finite-size scaling analysis at criticality of the first Lee-Yang zero from fits of the form $h_{0}L^{2.4819}\sim a(1+bL^{-0.832}+cL^{-1.168})$ [panel (a)] and $h_{0}L^{2.4819}\sim a(1+bL^{-0.832})$ panel (b)]. Similar analysis for the $8^{\rm th}$-order magnetization cumulant using now fits of the form $\langle\langle M^{8}\rangle\rangle L^{-16.8552}\sim a(1+bL^{-0.832}+cL^{-1.168})$ [panel (c)] and $\langle\langle M^{8}\rangle\rangle L^{-16.8552}\sim a(1+bL^{-0.832})$ [panel (d)].
@@ III.1.2 Accurate determination of the critical temperature
Figure 3: Finite-size scaling analysis of the density of the first Lee-Yang zero close to criticality for three values of the temperature, as indicated. Note that all three curves merge in the figure and are almost superimposed.
@@
In Fig. 3, we present the results for three different temperatures, as indicated in the panel. The curves are nearly indistinguishable, yet the best fit quality is achieved for the data corresponding to $T_{\rm c}=3.1962$, further reinforcing the findings from the previous analysis based on the first Lee-Yang zero and the magnetization cumulant. Additionally, the measured value for the exponent, $(d/y_{h})\approx 1.2057$ is in excellent agreement with the expected value, showing a deviation of less than $1\%$.
To move forward, it is now instructive to analyze more traditional quantities, namely the standard thermodynamic quantities. This will not only help us verify whether the results provide a consistent picture but also allow us to assess the advantages of our current approach. An estimate of the critical temperature is typically found by studying the crossings of the Binder cumulant (or of some other universal ratio, like $\xi/L$, where $\xi$ the correlation length) or via the shift behavior of related pseudocritical temperatures measured for the various diverging quantities (for example the location of the peaks of the specific heat or susceptibility). We choose to examine this latter route here for a set of thermodynamic observables defined above in Sec. II.3. As one also expects corrections to scaling, the scaling law describing the shift behavior of these pseudocritical temperatures $T_{L}$ reads as
@@ III.2 Tricritical scaling behavior
As previous estimations in the literature place the tricritical point near $(\Delta_{\rm t},T_{\rm t})=(2.8446,1.4182)$, we shall explore several values of these parameters in the vicinity of this point. The finite-size scaling analysis at both estimates proposed in Refs. [28, 27] for the location of the tricritical point gives a positive slope for the curves of the Lee-Yang zeros – $\ln(h_{0}L^{5/2})$ versus $\ln(\ln L)$ – where a negative slope is expected due to the $a(\ln L)^{-1/6}$ prediction in Eq. (22). A similar discrepancy appears for the nth-order magnetization cumulant for which a negative slope is measured, whereas one expects a positive slope, namely $\sim(\ln L)^{n/6}$ in Eq. (23). We therefore focus on three neighboring temperatures: $T=1.4181$, $1.4182$, and $T=1.4183$. Additionally, we examine four nearby values of the crystal field within the range $\Delta=2.8441-2.8444$. The numerical data in Fig. 6 indicate that the exponents are highly sensitive to the value of the crystal field, with variations observed even up to the fourth decimal place. Knowing that the expected exponent is $\hat{y}_{h}^{\rm(ref)}=1/6$, we find compatible results when $\Delta=2.8442$, along with a very good quality of fit. Indeed, at $T=1.4181$, the exponent value and the fit quality are slightly less accurate compared to those at $T=1.4182$ and $T=1.4183$. For these latter temperatures we obtain $\hat{y}_{h}=0.172(8)$ with $\chi^{2}_{s}=1.68$ and $\hat{y}_{h}=0.158(8)$ with $\chi^{2}_{s}=1.35$.
@@ Acknowledgements.
This work began in collaboration with our dear friend Ralph Kenna, who sadly passed away a year ago. In his memory, we would like to highlight two special issues published in tribute to him: one in the Ukrainian journal Condensed Matter Physics [51] published by the ICMP, with which Ralph had strong and fruitful collaborations, and another one in the journal Entropy [52], where he served on the editorial board. L.M. would like to express her gratitude to Ralph Kenna who was her PhD co-supervisor, for equipping her with the essential tools to advance her work. L.M. also thanks Coventry University, the L4 collaboration Leipzig-Lorraine-Lviv-Coventry and the UFA (French-German University) through the Collège Doctoral franco-allemand 02-07 for their support, and Denis Gessert for enriching discussions. N.G.F. would like to dedicate this paper to the memory of Ralph Kenna, an inspiring scientist and close colleague for more than 10 years at Coventry University. The work of N.G.F. was supported by the Engineering and Physical Sciences Research Council (grant EP/X026116/1 is acknowledged).